OpenStudy (trojanpoem):

Find the intervals of upwards concavity and downwards concavity and turning points if exists for the curve of function f as f(x) = x^2 +9 /x and find the absolute maximum and minimum for the function when x (- [1, 6]

OpenStudy (anonymous):

start by taking the first and second derivatives

OpenStudy (trojanpoem):

I know the steps, the problem lays in this example specifically. No turning points although there is upwards and downwards concavity.

OpenStudy (anonymous):

there is a vertical asymptote where denominator is equal to 0

OpenStudy (trojanpoem):

Clarify a bit .

OpenStudy (trojanpoem):

You mean something like this ? http://hotmath.com/hotmath_help/topics/rational-functions/image008.gif

OpenStudy (trojanpoem):

So there is no turning points as the curve is separated into two parts ?

OpenStudy (anonymous):

\[f(x)=\frac{x^2+9}{x}=x+\frac{9}{x}\] right?

OpenStudy (trojanpoem):

yep.

OpenStudy (trojanpoem):

I will watch the video while answering

OpenStudy (anonymous):

\[f'(x)=1-\frac{9}{x^2}\]find the interval over what \(f'\) is positive, that will tell you where it \(f\)is increasing

OpenStudy (anonymous):

i always graph these problems first to get an idea of what im looking to identify

OpenStudy (anonymous):

as for the absolute max on \([1,6]\) check the critical point (where \(f'(x)=0\) and also check \(f(1)\) and \(f(6)\)

OpenStudy (trojanpoem):

Good way of doing it, x = 0 -8, 0 > dec 0 , 8 increase 8 = infinite

OpenStudy (trojanpoem):

From now , I will graph it too.

OpenStudy (trojanpoem):

WHere is the turning point satellite ?

OpenStudy (anonymous):

in any case if you want a nice picture and pretty much everything else use this http://www.wolframalpha.com/input/?i=%28x^2%2B9%29%2Fx

OpenStudy (trojanpoem):

turning point satellite ?

OpenStudy (anonymous):

if "turning point" means where it goes from increasing to decreasing and vice versa, those are the zeros of the derivative

OpenStudy (trojanpoem):

Fine, so x= ?

OpenStudy (anonymous):

idk i didn't do it solve \[1-\frac{9}{x^2}=0\] and you will get them can probably do it in your head

OpenStudy (trojanpoem):

@billj5 , Finding the asymptote takes decade !

OpenStudy (trojanpoem):

This is the first derivative , we need the second one. and x = 0

OpenStudy (anonymous):

lol its not too bad once you get the hang of it

OpenStudy (trojanpoem):

At least , I understood why x = 0 was not a turning point :D , Thanks !

OpenStudy (trojanpoem):

Hey , may you help me with another question ? Will open new one.